# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ V.4 Properties in the vicinity of the critical temperature
A scaling law for the magnetic field valid for large fields and close to $T_{\mathrm{g}}$ was developed in Zhai et al. [300]. It was based on the assumption that a scaling theory that describes the magnetic response to an external field $H$ in equilibrium [238, 14] also holds in the non-equilibrium regime. Defining an “effective” response time, $t^{\text{eff}}_{H}$ in the same sense as Eq. (17) above, Eqs. (5558) can be written as
@@
The terms in Eq.  (64) are defined in the text following Eq. (55). Fits of Eqs. (57) and (58) to the data exhibit non-linearities in $H^{2}$, both from experiments and simulations, and are exhibited in Zhai et al. [300].
Figure 40: Effective time $t_{\mathrm{w}}^{\text{eff}}$ extracted from the time at which $S(t)$ peaks vs temperature for waiting times $t_{\mathrm{w}}$ from 300 to 100 000 s for a 6 at.CuMn single-crystal sample ($T_{\mathrm{g}}=31.5$ K). Reproduced from Kenning [171]
A spectacular example of the need for the scaling law is the collapse of $t^{\text{eff}}_{H}$ near $T_{\mathrm{g}}$ [171]. Fig. 40 exhibits his extraction of $t^{\text{eff}}_{H}$ for a CuMn 6 at.($T_{\mathrm{g}}=31.5$ K) as $T\rightarrow T_{\mathrm{g}}$ from below.
This collapse is evidence of the importance of the higher-order terms in the scaling law [Eqs. (57) and (58)]. The values of $a_{2}$ and $a_{4}$ were found by Zhai et al. [302] to be of opposite sign, so that the $H^{4}$ term is of opposite sign to the $H^{2}$ term in the expansion of Eq. (63). The power of the coherence length in the $H^{4}$ term is roughly twice the power of the coherence length in the $H^{2}$ term. Hence, as one approaches $T_{\mathrm{g}}$ [and $\xi_{\mathrm{micro}}$ increases, 34], the $H^{4}$ term rapidly grows compared to the $H^{2}$ term, sharply diminishing $t^{\text{eff}}_{H}$.
@@ V.5 Brief summary of dynamics in a field
The equilibrium and out-of-equilibrium dynamics of the 3D Ising spin glass in the presence of an external magnetic field was addressed in Baity-Jesi et al. [25] using numerical simulations with the Janus I dedicated supercomputer. The next paragraphs briefly review the main findings of this study.
The dynamics appear to be activated, in agreement with the droplet-model predictions that no phase transition occurs in a field. However the detailed dependence of the observables with time differs from that predicted by the droplet model. In particular, the data follows a super-Arrhenius law at temperatures as low as $T=0.36T_{\mathrm{g}}^{(H=0)}$. The large value of the correlation length at low temperatures indicates that the lower critical dimension is at most slightly greater than three.
Furthermore, equilibrium and out-of-equilibrium approaches allow the determination of a kind of transition line. However, the correlation length does not diverge on this line, against replica symmetry breaking (RSB) predictions. In addition, if the transition is driven by a $T=0$ fixed point (as in the droplet model) one should expect some kind of activated dynamics also in RSB. Thus, in this scenario, the identified transition line could be interpreted as a dynamic glass transition with the correlation length diverging at lower temperatures.
@@
In summary, the dynamical behavior in and out of equilibrium of the three-dimensional spin glass in a magnetic field exhibits large values of the correlation length. But this behavior is not compatible with the onset of a continuous phase transition (a de Almeida Thouless line). These facts, together with the results of equilibrium numerical simulations [26, 11], suggest that the lower critical dimension of the Ising spin-glass model in a field is greater than three but less than four [see also 232].
VI Off-equilibrium dynamics: non-linear response when varying the temperature
Rejuvenation and memory experiments [161] are among the most profound in the field of spin glasses. These experiments suggested that the study of magnetic responses in temperature-varying protocols could demonstrate some of the most interesting features of a complex free-energy landscape (most notably, the multiplicity of equilibrium states and temperature chaos). Once again, it is necessary to relate equilibrium landscapes to non-equilibrium experiments. This section reports on a continuing effort, originating in Bouchaud et al. [67] and still ongoing [31, 125, 230], to learn about the geometry of the free-energy landscape from temperature-varying experiments.
@@ VI.1 Temperature chaos in the off-equilibrium context
Temperature chaos was introduced as an equilibrium phenomenon in Secs. II.4 and III, while experiments are always conducted out of equilibrium. This section will explore how one can translate the concept of temperature chaos to the experimental setting and, then, how it can be investigated in non-equilibrium simulations.
@@ VI.1.1 The experimental point of view
Figure 41: The reference and the temperature cycling curves for $M_{\text{FC}}(t)$ of a CuMn 6 at.sample for $T_{1}=18$ K, and $t_{\mathrm{w1}}=10^{4}$ s. The temperature cycling curve is shifted by $\delta t$ to overlap the reference curve. In the reversible range, (a) and (b), the cycling curve can be overlapped with the reference curve over the whole period of measurement, $\approx 7\times 10^{4}$ s. In the chaotic range, (c) and (d), the cycling curve can only be partially overlapped. Hence, temperature chaos sets in for $T_{1}=18$ K, $t_{\mathrm{w1}}=10^{4}$ s at $\Delta T\geq 450$ mK. Reproduced from Fig. 1 of Zhai et al. [301].
@@
The search for temperature chaos in spin glasses has had a checkered history. The first systematic report [163] utilized both positive and negative temperature shifts for $\Delta T$. Arai et al. [18] were able to change bond strengths to access the chaotic state. Jönsson et al. [163] found for small $\Delta T$ that aging was “accumulative” but for larger $\Delta T$ a transition to “non-accumulative” aging was observed.
We associate accumulative to reversible behavior, and non-accumulative to chaos. Jönsson et al. [163] used ZFC measurements to extract the temperature ranges for accumulative and non-accumulative aging. They estimated the chaos exponent $(D_{\text{s}}/2-y)$ as
@@
The problem, therefore, is how to extract $1/(D_{\text{s}}/2-y)$ from experiment without changing a magnetic field. A solution was developed by Zhai et al. [301], who utilized the (small but finite) time dependence of the field-cooled magnetization, $M_{\text{FC}}(t)$. In this protocol, the magnetic field remains constant, so that only a change in temperature can contribute to chaos.
The protocol was first to establish a reference curve. A CuMn 6 at.single crystal ($T_{\mathrm{g}}=31.6$ K) was cooled from 40 K to the measurement temperature $T_{1}=18$ K in a constant field $H=40$ Oe at 10 K/min. After the temperature was stabilized, the reference magnetization $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ was recorded. In the temperature-cycling protocol, the sample was cooled from 40 K to $T_{1}=18$ K at 10 K/min, where the magnetization was recorded for a duration of $10^{4}$ s. The temperature was then lowered to $T_{2}=T_{1}-\Delta T$. After aging for $10^{3}$ s, the temperature was raised back to $T_{1}=18$ K, and the magnetization $M_{\text{FC}}^{\text{cyclic}}(t)$ recorded.
If there is no chaos, the two time-dependent magnetizations, $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ and $M_{\text{FC}}^{\text{cyclic}}(t)$, can be overlapped directly by shifting $M_{\text{FC}}^{\text{cyclic}}(t)$ by $\delta t$ [see Fig. 41, curves (a) and (b)], and aging can be taken as cumulative and reversible. However, if chaos is present, $M_{\text{FC}}^{\text{cyclic}}(t)$ cannot be made to overlap with $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ by shifting $M_{\text{FC}}^{\text{cyclic}}(t)$ by $\delta t$ [see Fig. 41, curves (c) and (d)].
Figure 41 displays the results of these experiments. Comparing Fig. 41 (b) with Fig. 41 (c) shows that temperature chaos begins between 17.55K and 17.54 K, or between $\Delta T=0.45$ K and 0.46 K, for an unprecedented accuracy of 10 mK. The amplitude of the lack of overlap increases with increasing $\Delta T$, as can be seen in Fig. 41 (d) where $\Delta T=12$ K. Zhai et al. [301] display additional values for $\Delta T$ for $T_{1}=18$ K, and for different $T_{1}=17-14$ K. Having these different $T_{1}$, and the concomitantly different $\Delta T$ for the onset of temperature chaos, enables the extraction of $1/(D_{\text{s}}/2-y)\approx 1.1$, well in the range of accepted values for $(D_{\text{s}}/2-y)$ from theoretical calculations and simulations for the 3D Ising model [see 301, for a list of these values].
@@ VI.1.2 The numerical point of view
The crucial elements to successfully observe temperature chaos in non-equilibrium simulations are three: the capacity to reach long times (enabled by specialized hardware, see Sec. II.7), the idea that temperature chaos is a rare event [111, see Sec. III], and the statics-dynamics equivalence [41, 43, 13, 36, see Secs. IV.3 and IV.4].
@@
Figure 42: Randomly chosen spheres of a sample of size $L=160$ are depicted with a color code depending on $1-X$ ($X$ is the chaotic parameter computed for spheres of radius $r=12$ and $\xi_{\mathrm{micro}}=12$ and temperature $T_{1}=0.7$ and $T_{2}=1.0$). For visualization purposes, spheres are represented with a radius $12(1-X)$, so that only fully chaotic spheres (i.e., $X=0$) have their real size. Figure from [30].
At this point, it is necessary to, somehow, compare configurations with different temperatures $T$ of the thermal bath. The natural choice is to use the chaotic parameter $X^{J}_{T_{1},T_{2}}$, see Eq. (13). While at equilibrium computing the chaotic parameter is straightforward, in off-equilibrium dynamics there is an extra parameter to play with: it is necessary to select the time $t$ of the compared configurations. The statics-dynamics equivalence suggests that, if in equilibrium the relevant length scale is the linear size $L$, the non-equilibrium system should be governed by the coherence length $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$, whose growth strongly depends on $T$ (see Sec. IV.2). Thus, the authors used the coherence length of the system as a proxy of its aging state and compared configurations at different times by keeping $\xi_{T_{1}}(t_{\mathrm{w1}})\approx\xi_{T_{2}}(t_{\mathrm{w2}})$. This strong condition will be discussed again in Sec. VI.2.2.
When studying the chaotic parameter as a global observable, the authors observed almost no signal. Again, the statics-dynamics equivalence suggests that the non-equilibrium system with $L\gg\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ can be considered as a collection of equilibrated systems with $L\sim\xi_{\mathrm{micro}}(t_{\mathrm{w}})$. Therefore, computing global observables implies averaging over equilibrated regions. As explained in Sec. III.1, this averaging washes the chaotic signal away.
Hence, the study of temperature chaos has to be local. The authors defined spherical regions of size $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ and computed within them a chaotic parameter
@@
where the superindex $\{s,r\}$ labels the specific sphere $s$ and the radius of the considered sphere $r$ for which we are computing the chaotic parameter and the thermal averages $\langle\cdots\rangle_{J}$ are as in Eq. (13). This local observable emphasizes the heterogeneity of the phenomenon, as shown in Fig. 42, in which there is a vast majority of non-chaotic spheres (blue-like colors) and specific regions with very chaotic spheres (red-like colors).
@@
A more quantitative study was conducted in Baity-Jesi et al. [30] by computing the probability distribution function of the chaotic parameter of the spheres,
@@
As sketched above, it was found that the chaotic signal peaks for a specific value of the radius of the studied sphere. The analysis of the peaks found two main results. First, the optimum radius to study temperature chaos is proportional to the coherence length $\xi_{\mathrm{micro}}$. Second, there exists a characteristic coherence length $\xi_{\mathrm{micro}}^{*}$ that points to a crossover between a weak-chaotic regime and a strong-chaotic regime. Moreover, since $\xi_{\mathrm{micro}}^{*}$ indicates the crossover between weak and strong chaos, it must be the non-equilibrium analog of the equilibrium chaotic length $\ell_{\text{c}}$ (see Sec. II.4). The authors found that the crossover length $\xi_{\mathrm{micro}}^{*}$ admits a scaling with the temperature difference $T_{2}-T_{1}$ of the form
@@ VI.2.1 An experimental approach
One of the remarkable features of spin-glass dynamics is their behavior under aging and temperature cycling, namely rejuvenation and memory. These properties were first exhibited by Jonason et al. [162] and are reproduced in Fig. 44. The imaginary part of the ac magnetic susceptibility, $\chi^{\prime\prime}(\omega)$, of the insulating spin glass CdCu1.7In0.3S4 ($T_{\mathrm{g}}=16.7$ K) as a function of temperature $T$ was first measured without aging, providing the reference curve (solid line in Fig. 44). Then the experiment was repeated, but this time as the temperature was lowered, it was stopped at $T=12$ K, where the system aged for 7 h. Then the temperature was lowered to the final $T=5$ K. Remarkably, $\chi^{\prime\prime}(\omega)$ returned to the reference curve after a small change in $T$. This behavior was termed “rejuvenation”: The measured $\chi^{\prime\prime}(\omega)$ behaved as if it had never been aged at $T=12$ K! This effect has been attributed to temperature chaos, but there remains need for a definitive proof of this conjecture.
@@
What was seen next was even more unexpected. As the system was heated from its lowest temperature, $\chi^{\prime\prime}(\omega)$ retraced its behavior, reproducing the drop seen during aging when the temperature was decreasing. This was termed “memory”. That it occurred after the system had rejuvenated makes it remarkable. For if rejuvenation is (and was) ascribed to a transition to a chaotic state, how could it “remember” what has occurred at a previous higher temperature after aging?
Figure 45: Measurements of the ac susceptibility displaying memory as a function of temperature. The imaginary part of the 1 Hz magnetic susceptibility, $\chi^{\prime\prime}(\omega)$. The dashed data are the reference curve measured while continuously cooling the sample at 1 K/min. the dot-dashed data include waiting at $T_{2}=22.5$ K for a time $t_{\mathrm{w1}}=1$ h and rejuvenation upon lowering the temperature. The temperature cooling rate is also 1 K/min. The solid curve is the heating data taken at a rate of 1 K/min that exhibits the memory effect at $T_{1}$. The inset shows an example of a double-waiting-time experiment. Reproduced from Fig. 1 of Freedberg et al. [125].
@@
Figure 46: ${\mathcal{M}}$ vs waiting time with four different first waiting temperatures. For both (a) and (b), the first and second waiting temperatures are the same: $T_{1}=18,~{}22.5,~{}27.5$, and 30 K and $T_{1}-T_{2}=$ 4 K. In both cases either the first (a) or second (b) waiting times were varied from 0 - 6 h (the horizontal axis), while the other waiting time was fixed at 3 and 1 h, respectively. Closed (open) markers indicate a variation of the first (second) waiting time. Reproduced from Fig. 2 of Freedberg et al. [125].
Figure 47: Memory vs waiting time data plotted against a model for memory loss predicted by [125]. Additionally, the three-starred data points use coherence lengths extracted from dc experiments with the ac values of memory. The data collapse supports the model. $T_{1}$ (18, 22.5, 27.5 and 30 K) and $T_{1}-T_{2}=4$ K. In both cases either the first (closed markers) or second (open markers) waiting times are varied from 0 to 6 h, while the other waiting time is fixed at 3 and 1 h, respectively. Statistical error bars are present on each data point, though they are typically smaller than the marker. Reproduced from Fig. 3 of Freedberg et al. [125].
@@
As to the length scale, their data show there must be another length scale at play other than the coherence length that is on the order of 100 nm. The origin of this length scale is not understand but may be related to the equilibrium correlation length.
@@
Figure 49: The ZFC numerical experiment measuring rejuvenation and memory. At time $t_{\mathrm{w}}$, a magnetic field $H=0.01$ is applied and the magnetic density, $M_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$, is recorded. Left panels show aging. The relaxation function $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ is displayed for the native runs at the warmer, $T_{1}=0.9$, and colder, $T_{2}=0.5$, temperatures . In the temperature-varying protocol (schematized by the green arrows), after a waiting time $t_{\mathrm{w1}}=2^{31.25}$ at the cold temperature $T_{1}=0.9$, the system is quenched to the cold one $T_{2}=0.5$ Then, the system relaxes at $T_{2}$ for an additional time, after which the magnetic field is switched on and the function $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ is measured (bottom-center). Finally, after the waiting time $t_{\mathrm{w}}^{\uparrow}=t_{\mathrm{w1}}+t_{\mathrm{w2}}=2^{32.25}$ (i.e., the system has spent half its life at the initial temperature $T_{1}$ and half at the colder temperature $T_{2}$ without a field), the spin glass is suddenly heated back to $T_{1}$. The system is allowed to relax for a short time, $t_{\mathrm{w}}=2^{10}\ll t_{\mathrm{w1}}$, after which the magnetic field is switched on. The $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ measured after the jump back (top-right) has a peak very similar to the one before the first jump (top-left), evincing the memory of the aging at the initial temperature $T_{1}$, notwithstanding the rejuvenation observed when staying at the lower temperature $T_{2}$. Figure taken from Baity-Jesi et al. [31].

### v2

@@ V.4 Properties in the vicinity of the critical temperature
A scaling law for the magnetic field valid for large fields and close to $T_{\mathrm{g}}$ was developed in Zhai et al. [375]. It was based on the assumption that a scaling theory that describes the magnetic response to an external field $H$ in equilibrium [296, 16] also holds in the nonequilibrium regime. Defining an “effective” response time, $t^{\text{eff}}_{H}$ in the same sense as Eq. (16) above, Eqs. (5962) can be written as
@@
The terms in Eq.  (68) are defined in the text following Eq. (59). Fits of Eqs. (61) and (62) to the data exhibit nonlinearities in $H^{2}$, both from experiments and simulations, and are exhibited in Zhai et al. [375].
Figure 41: Effective time $t_{\mathrm{w}}^{\text{eff}}$ extracted from the time at which $S(t)$ peaks vs temperature for waiting times $t_{\mathrm{w}}$ from 300 to 100 000 s for a 6 at.% CuMn single-crystal sample ($T_{\mathrm{g}}=31.5$ K). Reproduced from Kenning [205].
A spectacular example of the need for the scaling law is the collapse of $t^{\text{eff}}_{H}$ near $T_{\mathrm{g}}$ [205]. Fig. 41 exhibits his extraction of $t^{\text{eff}}_{H}$ for a CuMn 6 at.% ($T_{\mathrm{g}}=31.5$ K) as $T\rightarrow T_{\mathrm{g}}$ from below.
This collapse is evidence of the importance of the higher-order terms in the scaling law [Eqs. (61) and (62)]. The values of $a_{2}$ and $a_{4}$ were found by Zhai et al. [376] to be of opposite sign, so that the $H^{4}$ term is of opposite sign to the $H^{2}$ term in the expansion of Eq. (67). The power of the coherence length in the $H^{4}$ term is roughly twice the power of the coherence length in the $H^{2}$ term. Hence, as one approaches $T_{\mathrm{g}}$ [and $\xi_{\mathrm{micro}}$ increases, 36], the $H^{4}$ term rapidly grows compared to the $H^{2}$ term, sharply diminishing $t^{\text{eff}}_{H}$.
@@ V.5 Brief summary of dynamics in a field
The equilibrium and out-of-equilibrium dynamics of the 3D Ising spin glass in the presence of an external magnetic field were addressed in Baity-Jesi et al. [27] using numerical simulations with the Janus I dedicated supercomputer. The next paragraphs briefly review the main findings of this study.
The dynamics appear to be activated, in agreement with the droplet-model predictions that no phase transition occurs in a field. However, the detailed dependence of the observables with time differs from that predicted by the droplet model. In particular, the data follows a super-Arrhenius law at temperatures as low as $T=0.36T_{\mathrm{g}}^{(H=0)}$. The large value of the correlation length at low temperatures indicates that the lower critical dimension is at most slightly greater than three.
Furthermore, equilibrium and out-of-equilibrium approaches allow the determination of a kind of transition line. However, the correlation length does not diverge on this line, against replica-symmetry breaking (RSB) predictions. In addition, if the transition is driven by a $T=0$ fixed point (as in the droplet model) one should expect some kind of activated dynamics also in RSB. Thus, in this scenario, the identified transition line could be interpreted as a dynamic glass transition with the correlation length diverging at lower temperatures.
@@
In summary, the dynamical behavior in and out of equilibrium of the three-dimensional spin glass in a magnetic field exhibits large values of the correlation length. But this behavior is not compatible with the onset of a continuous phase transition (a de Almeida Thouless line). These facts, together with the results of equilibrium numerical simulations [28, 13], suggest that the lower critical dimension of the Ising spin-glass model in a field is greater than three but less than four [see also 289].
VI Off-equilibrium dynamics: nonlinear response when varying the temperature
Rejuvenation and memory experiments [190] are among the most profound in the field of spin glasses. These experiments suggested that the study of magnetic responses in temperature-varying protocols could demonstrate some of the most interesting features of a complex free-energy landscape (most notably, the multiplicity of equilibrium states and temperature chaos). Once again, it is necessary to relate equilibrium landscapes to nonequilibrium experiments. This section reports on a continuing effort, originating in Bouchaud et al. [72] and still ongoing [33, 147, 287], to learn about the geometry of the free-energy landscape from temperature-varying experiments.
@@ VI.1 Temperature chaos in the off-equilibrium context
Temperature chaos was introduced as an equilibrium phenomenon in Secs. II.4 and III, while experiments are always conducted out of equilibrium. This section will explore how one can translate the concept of temperature chaos to the experimental setting and, then, how it can be investigated in nonequilibrium simulations.
@@ VI.1.1 The experimental point of view
Figure 42: The reference and the temperature-cycling curves for $M_{\text{FC}}(t)$ of a CuMn 6 at.% sample for $T_{1}=18$ K, and $t_{\mathrm{w1}}=10^{4}$ s. The temperature-cycling curve is shifted by $\delta t$ to overlap the reference curve. In the reversible range, (a) and (b), the cycling curve can be overlapped with the reference curve over the whole period of measurement, $\approx 7\times 10^{4}$ s. In the chaotic range, (c) and (d), the cycling curve can only be partially overlapped. Hence, temperature chaos sets in for $T_{1}=18$ K, $t_{\mathrm{w1}}=10^{4}$ s at $\Delta T\geq 450$ mK. Reproduced from Fig. 1 of Zhai et al. [375].
@@
The search for temperature chaos in spin glasses has had a checkered history. The first systematic report [192] utilized both positive and negative temperature shifts for $\Delta T$. Arai et al. [20] were able to change bond strengths to access the chaotic state. Jönsson et al. [192] found for small $\Delta T$ that aging was “accumulative” but for larger $\Delta T$ a transition to “nonaccumulative” aging was observed.
We associate accumulative to reversible behavior, and nonaccumulative to chaos. Jönsson et al. [192] used ZFC measurements to extract the temperature ranges for accumulative and nonaccumulative aging. They estimated the chaos exponent $(D_{\text{s}}/2-y)$ as
@@
The problem, therefore, is how to extract $1/(D_{\text{s}}/2-y)$ from experiment without changing the magnetic field. A solution was developed by Zhai et al. [375], who utilized the (small but finite) time dependence of the field-cooled magnetization, $M_{\text{FC}}(t)$. In this protocol, the magnetic field remains constant, so that only a change in temperature can contribute to chaos.
The protocol was first to establish a reference curve. A CuMn 6 at.% single crystal ($T_{\mathrm{g}}=31.6$ K) was cooled from 40 K to the measurement temperature $T_{1}=18$ K in a constant field $H=40$ Oe at 10 K/min. After the temperature was stabilized, the reference magnetization $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ was recorded. In the temperature-cycling protocol, the sample was cooled from 40 K to $T_{1}=18$ K at 10 K/min, where the magnetization was recorded for a duration of $10^{4}$ s. The temperature was then lowered to $T_{2}=T_{1}-\Delta T$. After aging for $10^{3}$ s, the temperature was raised back to $T_{1}=18$ K, and the magnetization $M_{\text{FC}}^{\text{cyclic}}(t)$ recorded.
If there is no chaos, the two time-dependent magnetizations, $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ and $M_{\text{FC}}^{\text{cyclic}}(t)$, can be overlapped directly by shifting $M_{\text{FC}}^{\text{cyclic}}(t)$ by a time interval $\delta t$ [see Fig. 42, curves (a) and (b)], and aging can be taken as cumulative and reversible. However, if chaos is present, $M_{\text{FC}}^{\text{cyclic}}(t)$ cannot be made to overlap with $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ by shifting $M_{\text{FC}}^{\text{cyclic}}(t)$ by any $\delta t$ [see Fig. 42, curves (c) and (d)].
Figure 42 displays the results of these experiments. Comparing Fig. 42 (b) with Fig. 42 (c) shows that temperature chaos begins between 17.55K and 17.54 K, or between $\Delta T=0.45$ K and 0.46 K, for an unprecedented accuracy of 10 mK. The amplitude of the lack of overlap increases with increasing $\Delta T$, as can be seen in Fig. 42 (d) where $\Delta T=12$ K. Zhai et al. [375] display additional values for $\Delta T$ for $T_{1}=18$ K, and for different $T_{1}=17-14$ K. Having these different $T_{1}$, and the concomitantly different $\Delta T$ for the onset of temperature chaos, enables the extraction of $1/(D_{\text{s}}/2-y)\approx 1.1$, well within the range of accepted values for $(D_{\text{s}}/2-y)$ from theoretical calculations and simulations for the 3D Ising model [see 375, for a list of these values].
@@ VI.1.2 The numerical point of view
The crucial elements to successfully observe temperature chaos in nonequilibrium simulations are three: the capacity to reach long times (enabled by specialized hardware, see Sec. II.8), the idea that temperature chaos is a rare event [128, see Sec. III], and the statics-dynamics equivalence [43, 45, 15, 38, see Secs. IV.3 and IV.4].
@@
Figure 43: Randomly chosen spheres of a sample of size $L=160$ are depicted with a color code depending on 1X1-X ($X$ is the chaoticity parameter computed for spheres of radius $r=12$ and $\xi_{\mathrm{micro}}=12$ and temperature $T_{1}=0.7$ and $T_{2}=1.0$). For visualization purposes, spheres are represented with a radius 12(1X)12(1-X), so that only fully chaotic spheres (i.e., $X=0$) have their real size. Figure from [32].
At this point, it is necessary to, somehow, compare configurations with different temperatures $T$ of the thermal bath. The natural choice is to use the chaoticity parameter $X^{J}_{T_{1},T_{2}}$, see Eq. (10). While computing the chaoticity parameter at equilibrium is straightforward, in off-equilibrium dynamics there is an extra parameter to play with: it is necessary to select the time $t$ of the compared configurations. The statics-dynamics equivalence suggests that, if in equilibrium the relevant length scale is the linear size $L$, the nonequilibrium system should be governed by the coherence length $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$, whose growth strongly depends on $T$ (see Sec. IV.2). Thus, the authors used the coherence length of the system as a proxy of its aging state and compared configurations at different times by keeping $\xi_{T_{1}}(t_{\mathrm{w1}})\approx\xi_{T_{2}}(t_{\mathrm{w2}})$. This strong condition will be discussed again in Sec. VI.2.2.
When studying the chaoticity parameter as a global observable, the authors observed almost no signal. Again, the statics-dynamics equivalence suggests that the nonequilibrium system with $L\gg\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ can be considered as a collection of equilibrated systems with $L\sim\xi_{\mathrm{micro}}(t_{\mathrm{w}})$. Therefore, computing global observables implies averaging over equilibrated regions. As explained in Sec. III.1, this averaging washes the chaotic signal away.
Hence, the study of temperature chaos has to be local. The authors defined spherical regions of size $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ and computed within them a chaoticity parameter
@@
where the superindex $\{s,r\}$ labels the specific sphere $s$ and the radius of the considered sphere $r$ for which we are computing the chaoticity parameter and the thermal averages $\langle\cdots\rangle_{J}$ are as in Eq. (10). This local observable emphasizes the heterogeneity of the phenomenon, as shown in Fig. 43, in which there is a vast majority of nonchaotic spheres (blue-like colors) and specific regions with very chaotic spheres (red-like colors).
@@
A more quantitative study was conducted in Baity-Jesi et al. [32] by computing the probability distribution function of the chaoticity parameter of the spheres,
@@
As sketched above, it was found that the chaotic signal peaks for a specific value of the radius of the studied sphere. The analysis of the peaks found two main results. First, the optimum radius to study temperature chaos is proportional to the coherence length $\xi_{\mathrm{micro}}$. Second, there exists a characteristic coherence length $\xi_{\mathrm{micro}}^{*}$ that points to a crossover between a weakly chaotic regime and a strongly chaotic regime. Moreover, since $\xi_{\mathrm{micro}}^{*}$ indicates the crossover between weak and strong chaos, it must be the nonequilibrium analog of the equilibrium chaotic length $\ell_{\text{c}}$ (see Sec. II.4). The authors found that the crossover length $\xi_{\mathrm{micro}}^{*}$ admits a scaling with the temperature difference $T_{2}-T_{1}$ of the form
@@ VI.2.1 An experimental approach
One of the remarkable features of spin-glass dynamics is their behavior under aging and temperature cycling, namely rejuvenation and memory. These properties were first exhibited by Jonason et al. [191] and are reproduced in Fig. 45. The imaginary part of the ac magnetic susceptibility, $\chi^{\prime\prime}(\omega)$, of the insulating spin glass CdCu1.7In0.3S4 ($T_{\mathrm{g}}=16.7$ K) as a function of temperature $T$ was first measured without aging, providing the reference curve (solid line in Fig. 45). Then the experiment was repeated, but this time as the temperature was lowered, it was stopped at $T=12$ K, where the system aged for 7 h. Then the temperature was lowered to the final $T=5$ K. Remarkably, $\chi^{\prime\prime}(\omega)$ returned to the reference curve after a small change in $T$. This behavior was termed “rejuvenation”: The measured $\chi^{\prime\prime}(\omega)$ behaved as if it had never been aged at $T=12$ K! This effect was attributed to temperature chaos, but there remained need for a definitive proof of this conjecture.
@@
What was seen next was even more unexpected. As the system was heated from its lowest temperature, $\chi^{\prime\prime}(\omega)$ retraced its behavior, reproducing the drop seen during aging when the temperature was decreasing. This was termed “memory”. That it occurred after the system had rejuvenated makes it remarkable. For if rejuvenation is (and was) ascribed to a transition to a chaotic state, how could it “remember” what had occurred at a previous higher temperature after aging?
Figure 46: Measurements of the ac susceptibility displaying memory as a function of temperature. The imaginary part of the 1 Hz magnetic susceptibility, $\chi^{\prime\prime}(\omega)$. The dashed data is the reference curve measured while continuously cooling the sample at 1 K/min. The dot-dashed data includes waiting at $T_{2}=22.5$ K for a time $t_{\mathrm{w1}}=1$ h and rejuvenation upon lowering the temperature. The temperature-cooling rate is also 1 K/min. The solid curve is the heating data taken at a rate of 1 K/min that exhibits the memory effect at $T_{1}$. The inset shows an example of a double-waiting-time experiment. Reproduced from Fig. 1 of Freedberg et al. [147].
@@
Figure 47: $\mathcal{M}$} vs waiting time with four different first waiting temperatures. For both (a) and (b), the first and second waiting temperatures are the same: $T_{1}=18,~22.5,~27.5$, and 30 K and $T_{1}-T_{2}=$ 4 K. In both cases either the first (a) or second (b) waiting times were varied from 0 - 6 h (the horizontal axis), while the other waiting time was fixed at 3 and 1 h, respectively. Closed (open) markers indicate a variation of the first (second) waiting time. The lines are fits to their phenomenological model with the same set of three adjustable parameters. Reproduced from Fig. 2 of Freedberg et al. [147].
Figure 48: Memory vs waiting time data plotted against a model for memory loss predicted by [147]. Additionally, the three starred data points use coherence lengths extracted from dc experiments with the ac values of memory. The data collapse supports the model. $T_{1}$ (18, 22.5, 27.5 and 30 K) and $T_{1}-T_{2}=4$ K. In both cases either the first (closed markers) or second (open markers) waiting times are varied from 0 to 6 h, while the other waiting time is fixed at 3 and 1 h, respectively. Statistical error bars are present on each data point, though they are typically smaller than the marker. Reproduced from Fig. 3 of Freedberg et al. [147].
@@
As to the length scale, their data show there must be another length scale at play other than the coherence length that is on the order of 100 nm. The origin of this length scale is not understood but may be related to the equilibrium correlation length.
@@
Figure 50: The ZFC numerical experiment measuring rejuvenation and memory. At time $t_{\mathrm{w}}$, a magnetic field $H=0.01$ is applied and the magnetic density, $M_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$, is recorded. Left panels show aging. The relaxation function $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ is displayed for the native runs at the warmer, $T_{1}=0.9$, and colder, $T_{2}=0.5$, temperatures . In the temperature-varying protocol (schematized by the green arrows), after a waiting time $t_{\mathrm{w1}}=2^{31.25}$ at the hot temperature $T_{1}=0.9$, the system is quenched to the cold one $T_{2}=0.5$ Then, the system relaxes at $T_{2}$ for an additional time, after which the magnetic field is switched on and the function $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ is measured (bottom-center). Finally, after the waiting time $t_{\mathrm{w}}^{\uparrow}=t_{\mathrm{w1}}+t_{\mathrm{w2}}=2^{32.25}$ (i.e., the system has spent half its life at the initial temperature $T_{1}$ and half at the colder temperature $T_{2}$ without a field), the spin glass is suddenly heated back to $T_{1}$. The system is allowed to relax for a short time, $t_{\mathrm{w}}=2^{10}\ll t_{\mathrm{w1}}$, after which the magnetic field is switched on. The $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ measured after the jump back (top-right) has a peak very similar to the one before the first jump (top-left), evincing the memory of the aging at the initial temperature $T_{1}$, notwithstanding the rejuvenation observed when staying at the lower temperature $T_{2}$. Figure taken from Baity-Jesi et al. [33].